Swimming space
The length of a rectangular swimming pool is 2x – 1 meters, and the width is x + 2 meters. Write a polynomial that represents the area. Find the area if x is 5 meters.

i have a hard time with word problems

Answer by stanbon(75887)

The length of a rectangular swimming pool is 2x – 1 meters, and the width is x + 2 meters. Write a polynomial that represents the area. Find the area if x is 5 meters.
-------------
Area = length * width
A(x) = (2x-1)(x+2)
A(5) = (10-1)(5+2)=9*7=63 sq. meters
Cheers
Stan H.